OpenStudy (anonymous):

What is the period of y = -4 sin x/2? a. -4 b. 2 c. 4π d. π e. 4 f. 1/2

jimthompson5910 (jim_thompson5910):

the period of sin(Bx) is T = 2pi/B in this case, B = 1/2, so T = 2pi/(1/2) T = 2*2pi T = 4pi

OpenStudy (anonymous):

whats the amplitude please? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

The amplitude of y = A*sin(x) is |A|

jimthompson5910 (jim_thompson5910):

in this case, A = -4

OpenStudy (anonymous):

thats what i had put! thanks :)